OpenStudy (anonymous):

help please find the indicated sum for the arithmetic series problem below ...

OpenStudy (anonymous):

\[\sum_{k=1}^{12} (-2+6k)\]

OpenStudy (anonymous):

a) 444 b) 368 c) 888 d)565

OpenStudy (campbell_st):

find the 1st term when k = 1 and the last term when k = 12 1st term a = -2+6 * (1) = 4 last term l = -2 + 6*(12) = 70 the sum of the terms is \[S_{n} = \frac{n}{2}(a + l)\] you found a and l above, and n = 12 substitute and evaluate.

OpenStudy (anonymous):

its 444

OpenStudy (anonymous):

how'd you get that @breeza

OpenStudy (anonymous):

where would you substitute the number ? @campbell_st

OpenStudy (anonymous):

because on that bottom equation, he's telling you n=12 and a= the first term when u substitute 1 so a=4 and I= the last term you substituted and they wan you 2 substitute 12 so I=70

OpenStudy (anonymous):

juist plug in the numbers he gave you

OpenStudy (campbell_st):

if you read the information and plug in the values its \[S_{12} = \frac{12}{2}[ 4 + 70]\]